OpenStudy (anonymous):

3x+y=-21 and x+y=-5 solve system of equations

OpenStudy (ranga):

Subtract the second equation from the first to eliminate y. Solve for x.

OpenStudy (anonymous):

What's answer What's the answer

OpenStudy (anthonyn2121):

Solve for y for the second equation: y=-x-5. Substitute y into the first equation 3x+(-x-5)=-21. Combine like terms and solve for x. 2x-5=-21. x=-8

OpenStudy (anonymous):

THANK YOU!!!